# Parametric Instabilities of Correlated Quantum Matter

### v1

@@ IV.2 Complex orders in a flat band model
IV.2.1 Scenario T ($\Delta_{\sigma}=0$)
In the absence of the $\Delta_{\sigma}$ term, an appropriate order parameter is ${\cal O}_{\rm T}=\sigma_{x}\tau_{x}$, where the $\tau_{x}$ part has been arbitrarily chosen, as the valley $U\left(1\right)$ symmetry is spontaneously broken. The hierarchy of ordered ground states within the low-energy manifold we consider here is depicted in Fig. 4a. Restricting to collective fluctuations which leave the system in its low-energy manifold determined by Eq. (56555These are operators which generate a rotated reference state $|\Psi_{\phi}\rangle$ whose corresponding $\cal O$ matrix lies within the low-energy manifold., there are four relevant generators to consider, $\left\{{{\cal M}^{i}}\right\}=\tau_{z},\,\sigma_{x}\tau_{y},\,\sigma_{z},\,\sigma_{y}\tau_{x}$.
Consulting again with Eq. (15), we can pair $\tau_{z}$ with $\sigma_{x}\tau_{y}$ as conjugate variables, as well as $\sigma_{z}$ with $\sigma_{y}\tau_{x}$. One expects the first pair to give rise to a gapless Goldstone mode. This is because a $\tau_{z}$ rotation applied to the order parameter generates a very soft fluctuation, namely rotation of the order parameter in the $\tau_{x}-\tau_{y}$ plane. Its partner $\sigma_{x}\tau_{y}$ acting on the order parameter generates energetically-unfavorable valley-polarization. The difference in energetics of these two fluctuation quadratures leads to effective squeezing of this bosonic mode.
The second pair is squeezed as well, see Fig. 4b: $\sigma_{y}\tau_{x}$ generates $\sigma_{z}$ polarization, which (by our assumption of $u_{\zeta}>\Delta_{\rm TK}$) is disfavored compared to the K order generated by $\sigma_{z}$. Moreover, since the K-order requires finite energy compared to the ${\cal O}_{\rm T}$ reference state, the collective mode spectrum associated with this second pair of generators is expected to have a finite gap. The dispersion of these two low-lying modes is presented in Fig. 1b.   [^6] Interestingly, in this model the clear separation to two modes associated with these generator pairs is spoiled at higher momenta. Specifically, the interaction terms include finite coupling between the two modes at nonzero $\bf Q$, see Appendix F. Parametric driving of this coupling between collective modes of ordered electrons is a promising direction we leave for future work.
We find that at large wavelengths the gapless Goldstone mode has the effective parameters (Appendix F),
@@ IV.2.1 Scenario T ($\Delta_{\sigma}=0$)
The linear Goldstone dispersion is approximately ${\cal E}^{\rm Goldstone}_{\bf Q}\approx\sqrt{\rho_{s}\left(u_{\zeta}+\Delta_{\rm TK}\right)}\left|{\bf Q}\right|$. The gapped mode however, accounting for fluctuations into the disfavored K-manifold, is described by
@@
Notice that in the limit where $u_{\zeta}\to\Delta_{\rm TK}$, the mode is absent squeezing, confirming our expectations above. The energy gap in this mode is
@@
$\Delta_{\rm TK}\to\Delta_{\rm TK}+\delta\Delta_{\rm TK}\cos\left(\omega_{d}t\right).$
@@
Figure 4: Driving in the T-order phase of the flat band model (Sec. IV.2.1). (a) Energy hierarchy within the low-energy manifold of the possible ordered ground states. The energy $u_{\zeta}$ separates the T/K orders from the rest of the manifold, and $\Delta_{\rm TK}$ further favors T-order over K. (b) The generalized Bloch sphere for the T-order case, positioned the north pole. The two fluctuation quadratures correspond to the easy K-order, and the hard sublattice ($\sigma$) order. The anisotropy of the fluctuation cloud corresponds to the squeezing in the fluctuation vacuum. Next to each order we schematically draw real-space spectroscopic signatures of the analogous TBG phase: T-order admits a Kekulé distortion of the graphene unit cell, $\sigma$ corresponds to finite sub-lattice polarization, and K-order is not discernible in such experiments  [118]. Modulating the parameter $\Delta_{\rm TK}$ affects the squeezing, and thus enables parametric resonance. In the corresponding driven state one observes substantial sublattice oscillations at half the driving frequency.
Since $\Delta_{\rm TK}$ originates in electrons interacting with a specific Raman-active phonon modes [103, 119, 120], driving a proxy IR-active phonon which couples to this phonon non-linearly is a route towards $\Delta_{\rm TK}$ modulation. Crucially, the envelope of this drive is externally set to oscillate at the parametric frequency $2{\cal E}^{\rm K}_{\bf 0}$ (and not the much higher phonon frequency itself). This induces time-dependency for the phonon parameters, and as a consequence modulates the phonon-mediated electronic interactions (cf. Ref. [121]).
Quantitatively the parametric susceptibility for this driving is readily extracted with the help of Eq. (38),
$\frac{\delta z^{\rm K}}{\delta\Delta_{\rm TK}}=-\frac{u_{\zeta}}{\sqrt{2\Delta_{{\rm TK}}\left(u_{\zeta}+\Delta_{{\rm TK}}\right)}}=-\frac{u_{\zeta}}{{\cal E}^{\rm K}_{\bf 0}}.$ (64)
@@
Two immediate consequences are apparent from Eq. (64). First, the susceptibility diverges as $\Delta_{\rm TK}\to 0$. This can be appreciated from the point of view of the fidelity susceptibility, Eq. (34). When $\Delta_{\rm TK}$ approaches zero, finally changing sign and crossing over, the “T” reference state is destabilized in favor of “K-order”. The vacuum of quantum fluctuations becomes increasingly sensitive to the transition, with $\chi^{F}\propto\Delta_{\rm TK}^{-2}$ in its vicinity. Second, probing the strength of the parametric response, one gains access not only to the gap of this mode (which is may be recovered independently by locating the resonance frequency), but also to secondary features of the ground state energy hierarchy via the quantity $u_{\zeta}$.
In fact, Eq. (64) does not tell the entire story in terms of the parametric efficiencies. The more physical question pertains to the sensitivity to different interaction terms (which are generally independent) captured by $H_{\rm K}$ and $H_{\rm T}$. Concretely,
$\eta_{{\rm par}.}^{\left(u^{{\rm T/K}}\right)}=\frac{1}{1+\frac{\Delta_{{\rm TK}}}{u_{\zeta}}}\frac{u^{{\rm T/K}}}{\Delta_{{\rm TK}}},$ (65)
which clearly exhibits the divergence near $\Delta_{\rm TK}\to 0$, where the phase transition between T-order and K-order takes place.
@@
What are the consequences of the resonant non-equilibrium state? The two pumped quadratures correspond to $\langle\sigma_{y}\tau_{x}\rangle$ – the easy K-order direction, and $\langle\sigma_{z}\rangle$ – the hard direction. (The easy quadrature has a relative amplitude of $r^{-1}=\frac{{\cal E}_{{\bf 0}}^{{\rm K}}}{2\Delta_{{\rm TK}}}$ compared to the harder one.) Remarkably, a phase which is otherwise inaccessible in an equilibrium setting, the K-IVC analog in our model, plays a key role in the dynamics of the driven state. Moreover, this state hosts an analog of an oscillating staggered sublattice potential. The resultant phase of Floquet matter, enabled by coherent manipulation of this strongly interacting system, potentially gives rise to unusual non-linear optical responses (e.g., second-harmonic generation [122, 123] due to the time-dependent inversion symmetry breaking) activated by pumping the system at a narrow frequency window.
IV.2.2 Scenario $\sigma$ ($\Delta_{\sigma}>\frac{u_{\zeta}+\Delta_{\rm TK}}{2}$)
@@
If the externally induced $\Delta_{\sigma}$ term (which in the case of graphene-based materials typically corresponds to alignment with a substrate which explicitly breaks the inversion/ sublattice symmetry) overwhelms the other interaction scales, the preferred order parameter for the system is ${\cal O}_{\sigma}=\sigma_{z}$. The relevant ordered state hierarchy for this case is shown in Fig. 5a. The generators which keep the system in the low energy manifold of Eq. (56) are thus $\left\{{\cal M}_{i}\right\}=\sigma_{x}\tau_{x},\,\sigma_{y}\tau_{x},\,\sigma_{x}\tau_{y},\,\sigma_{y}\tau_{y}$.
We may again organize then in canonically conjugate pairs, $\sigma_{x}\tau_{x}$ with $\sigma_{y}\tau_{x}$, and $\sigma_{x}\tau_{y}$ with $\sigma_{y}\tau_{y}$. The generator containing $\sigma_{y}$ in each pair generates a small amount of T-order in the rotated state, whereas the $\sigma_{x}$ generator creates K-order (see Fig. 5b). It is thus natural to expect that the corresponding collective modes will be squeezed only if $\Delta_{\rm TK}\neq 0$. The lowest-lying modes corresponding to the two generator pairs are further expected to be degenerate ($\tau_{x}$ and $\tau_{y}$ are on equal footing) and gapped (no continuous symmetry of the Hamiltonian is spontaneously broken by the ordered state).
@@
We note here that the ordering scenario described here is analogous to the antiferromagnet discussed in Fig. 2c, with $\Delta_{\rm\sigma}$ playing the role of a staggered external magnetic field. To see that this is the case, simply perform a Pauli matrix replacement which conserves all commutation and anti commutation relations ($s^{i}$ and $p^{i}$ are Pauli matrices describing fictitious “spin” and “sublattice” degrees of freedom): The antiferromagnetic order is now $\sigma_{z}\to s^{z}p^{z}$, Néel vector rotations (or T-order generators) are $\sigma_{y}\tau_{x}\to s^{x}$, $\sigma_{y}\tau_{y}\to s^{y}$, and ferromagnetic ordering (or K-order generators) is given by generators $\sigma_{x}\tau_{x}\to s^{x}p^{z}$, $\sigma_{x}\tau_{y}\to s^{y}p^{z}$.
@@
Indeed, as expected, the squeezing of lowest-lying collective modes is given by the breaking of T/K symmetry in the model. The modes have the expected gap as well at ${\bf Q}=0$, whose size is ${\cal E}^{\left(1,2\right)}_{\bf 0}=\sqrt{\left(\Delta_{\sigma}-\frac{u_{\zeta}}{2}\right)^{2}-\left(\frac{\Delta_{{\rm TK}}}{2}\right)^{2}}$. Unsurprisingly, the gap closes when the $\sigma_{z}$ order becomes exactly degenerate with the T-ordered state, i.e., when $\Delta_{\sigma}=\frac{u_{\zeta}+\Delta_{\rm TK}}{2}$. Let us now see how the vicinity to a phase transition affects the parametric susceptibility once again.
As an example, consider small changes in the Hamiltonian parameters that affect the quantum geometry of electrons in the projected bands, modifying slightly the quantum metric length $\zeta$. In realistic devices, this can be the result of modulating gate voltages, electric displacement fields, or strain. More exotic possibilities include modulation of external driving in Floquet systems, where the Floquet driving subtly controls single-particle properties [124]. The parametric efficiency to such a modification may be extracted with the help of Eq. (38),
@@
Notice that the left term in the denominator vanishes at the phase transition to the T-order, causing a divergence in the parametric efficiency (also see Fig. 5c). Once more, this can be understood from the fidelity susceptibility diverging at the transition.
The dependence on $\zeta$ comes from the derivative of $u_{\zeta}$, and is maximized for $\zeta\approx 0.5$. This particular value is not universal – the maximizing value will typically depend on the range of interactions. Under our present assumptions, the $u$ interactions are maximally short-ranged, thus a change of length scale associated with the projected band electrons, i.e., $\sim\zeta$, beyond a certain value has limited effect. Generalizing this observation, small changes in quantum geometrical parameters are expected to have the most pronounce impact when the length-scale associated with the interaction is comparable to the one derived from the quantum metric. The interplay between these energy scales has been demonstrated to be crucial in the context of fractional Chern insulators [125, 126]. A parametric response probe may thus shed much needed light on the prospect of realizing these phases in possible candidate materials.
Figure 5: Pseudospin (sublattice) polarized order and its parametric driving (Sec. IV.2.2). (a) Energy hierarchy with a finite external sublattice potential $\Delta_{\sigma}$ in Eq. (52a)]. The energy $u_{\zeta}$ brings the T/K orders down in energy, and $\Delta_{\rm TK}$ once again favors T-order over K. (b) The generalized Bloch sphere for the $\sigma$-order case, positioned the north pole. The two fluctuation quadratures correspond to the easy T-order, and the hard K-order, distinguished by the interaction energy $\Delta_{\rm TK}$. The anisotropy of the fluctuation cloud corresponds again to the vacuum. (c) Left: Parametric efficiency with respect to the quantum geometric parameter $\zeta$, Eq. (68). This parameter roughly corresponds to the Wannier spread in real-space of the wavefunctions corresponding to the active flat-bands (right schematic). The dashed line marks the phase boundary between the $\sigma$ phase below and the T-order stabilized above (hatched region), in accordance with the hierarchy in (a). Red line traces where $\eta_{{\rm par}.}^{\left(\zeta\right)}=1$. The parametric response is maximal when the interaction energy strongly depends on quantum geometry. The susceptibility diverges as one approaches the phase boundary, even at $\zeta$ much larger than its optimal value. Right: A vertical cut of the left colormap at $\zeta=0.65$, showing the divergence as the phase boundary is approached. Here, $\Delta_{\sigma}=u$.
@@
The response to driving in this scenario would be an oscillatory appearance of T-order and K-order components in the transient order parameter, with the amplitude of the former being stronger by the squeezing factor of $r^{-1}=\sqrt{1+\frac{\Delta_{{\rm TK}}}{\Delta_{\sigma}-\frac{u_{\zeta}+\Delta_{{\rm TK}}}{2}}}$ compared to the latter. We note that in the analogous twisted graphene system, the T-order, i.e., TIVC, couples to a real-space charge density with a Kekulé pattern [118, 128, 113, 103]. Thus, the parametrically driven state will give rise to a charge density wave oscillating at half the drive frequency, potentially detectable using available time-resolved STM techniques [129, 130].

### v2

@@ IV.2 Complex orders in a flat band model
IV.2.1 T Intervalley Coherence (Fig. 4b)
Without explicitly breaking the inversion symmetry, i.e., setting
$\Delta_{\sigma}=0,$
the order parameter is ${\cal O}_{\rm T}=\sigma_{x}\tau_{x}$, where the $\tau_{x}$ part has been arbitrarily chosen, as the valley $U\left(1\right)$ symmetry is spontaneously broken. The hierarchy of ordered ground states within the low-energy manifold we consider here is depicted in Fig. 6a. Restricting to collective fluctuations which leave the system in its low-energy manifold determined by Eq. (55555These are operators which generate a rotated reference state $|\Psi_{\phi}\rangle$ whose corresponding $\cal O$ matrix lies within the low-energy manifold., there are four relevant generators to consider, $\left\{{{\cal M}^{i}}\right\}=\tau_{z},\,\sigma_{x}\tau_{y},\,\sigma_{z},\,\sigma_{y}\tau_{x}$.
Consulting again with Eq. (16), we can pair the fluctuation amplitudes associated with the operators $\tau_{z}$ and $\sigma_{x}\tau_{y}$ as conjugate variables. The amplitudes corresponding to $\sigma_{z}$ and $\sigma_{y}\tau_{x}$ are thus a second conjugate pair. One expects the first pair to give rise to a gapless Goldstone mode. This is because a $\tau_{z}$ rotation applied to the order parameter generates a very soft fluctuation, namely rotation of the order parameter in the $\tau_{x}$-$\tau_{y}$ plane. Its partner $\sigma_{x}\tau_{y}$ acting on the order parameter generates energetically-unfavorable valley-polarization. The difference in energetics of these two fluctuation quadratures leads to effective squeezing of this bosonic mode.
The second pair of fluctuation matrices gives rise to squeezed fluctuations as well, see Fig. 6b. To understand why, consider that the $\sigma_{y}\tau_{x}$ piece of $\phi_{\bf Q}$ generates finite sub-lattice polarization in $|\Psi_{\phi}\rangle$. By our assumptions, this fluctuation is energetically disfavored compared to the KIVC component, whose generation stems from the $\sigma_{z}$ part of the fluctuation matrix $\phi_{\bf Q}$. Moreover, since the KIVC order requires finite energy compared to the TIVC ${\cal O}_{\rm T}$ reference state, the collective mode spectrum associated with this second pair of generators has a finite gap. The dispersion of these two low-lying modes is numerically calculated and presented in Fig. 2b.   [^6] Interestingly, in this model the clear separation to two modes associated with these generator pairs is spoiled at higher momenta. Specifically, the interaction terms include finite coupling between the two modes at nonzero $\bf Q$, see Appendix F. Parametric driving of this coupling between collective modes of ordered electrons is a promising direction we leave for future work.
At large wavelengths we can make significant analytical progress. The gapless Goldstone mode, comprised of $\tau_{z}$ and $\sigma_{x}\tau_{y}$ perturbations, has effective single-mode parameters (Appendix F),
@@
These parameters correspond to the Hamiltonian of the form of Eq. (37). The linear Goldstone dispersion is approximately ${\cal E}^{\rm Goldstone}_{\bf Q}\approx\sqrt{\rho_{s}\left(u_{\zeta}+\Delta_{\rm TK}\right)}\left|{\bf Q}\right|$.
The gapped mode however, whose generators are $\sigma_{z}$ and $\sigma_{y}\tau_{x}$, is similarly described by
@@
Notice that in the limit where $u_{\zeta}\to\Delta_{\rm TK}$, the mode has no squeezing ($\upsilon^{\rm K}_{\bf Q}\to 0$), confirming our expectations above – The KIVC and sublattice-polarized orders have identical Hartree-Fock energies. The energy gap for this mode is
@@
$\Delta_{\rm TK}\to\Delta_{\rm TK}+\delta\Delta_{\rm TK}\cos\left(2{\cal E}^{\rm K}_{\bf 0}t\right).$
@@
Figure 6: Fluctuations in the TIVC phase of the flat band model (Sec. IV.2.1). (a) Energy hierarchy within the low-energy manifold of the possible ordered ground states. The energy $u_{\zeta}$ separates the TIVC ($\sigma_{x}\tau_{x}$) and KIVC ($\sigma_{y}\tau_{x}$) orders from the rest of the manifold, and $\Delta_{\rm TK}$ further favors the TIVC. (b) The generalized Bloch sphere for the collective mode described by Eqs. (60)-(61), with the TIVC order parameter positioned at the north pole. The two fluctuation quadratures correspond to the easy KIVC, and the hard sublattice ($\sigma$) orders. The anisotropy of the schematic fluctuation cloud corresponds to the squeezing in the fluctuation vacuum. Next to each order we schematically draw real-space spectroscopic signatures of the analogous TBG phase: TIVC admits a Kekulé distortion of the graphene unit cell, $\sigma$ corresponds to finite sublattice polarization, and KIVC is not discernible in such experiments  [123]. Modulating the parameter $\Delta_{\rm TK}$ affects the squeezing, and thus enables parametric resonance. In the corresponding driven state one observes substantial sublattice oscillations at half the driving frequency.
Microscopically, $\Delta_{\rm TK}$ originates in electrons interacting with specific Raman-active phonon modes [82, 124, 125]. A promising route towards time-modulating this interaction is to optically pump an auxiliary IR-active phonon, which couples to one of these phonons non-linearly  [^7] See recent works on non-linear phononics based parametric resonances, Refs. [62, 168]. The amplitude of the pump drive should be externally such that the phonon-induced electronic interaction strength oscillates with frequency $2{\cal E}^{\rm K}_{\bf 0}$. With such experimental realization in mind, one expects that the direct effect of the driving would be to soften the interaction-inducing phonon modes (cf. Ref. [127]). The induced electronic interaction are approximately inversely proportional to the phonon frequency [82]. Thus, the relative amplitude $\delta\Delta_{\rm TK}/\Delta_{\rm TK}$ roughly corresponds to the relative softening of the characteristic frequency of the phonon modes $\omega_{\rm ph}$, i.e., to $\delta\omega_{\rm ph}/\omega_{\rm ph}$.
Thus, physically the modulation of $\Delta_{\rm TK}$ is coming mainly from the modification of the $u^{\rm T}$ interaction [Eq. (57)], which favors the phonon-driven TIVC phase. The parametric efficiency with respect to these interactions,
$\eta_{{\rm par}.}^{\left(u^{{\rm T}}\right)}=\frac{1}{1+\frac{\Delta_{{\rm TK}}}{u_{\zeta}}}\frac{u^{{\rm T}}}{\Delta_{{\rm TK}}}.$ (62)
@@
Notice that the parametric efficiency diverges as $\Delta_{\rm TK}\to 0$. This can be appreciated from the point of view of the fidelity susceptibility, Eq. (36). When $\Delta_{\rm TK}$ approaches zero, finally changing sign and crossing over, the TIVC reference state is destabilized in favor of KIVC. The vacuum of quantum fluctuations becomes increasingly sensitive to the transition, with $\chi^{F}\propto\Delta_{\rm TK}^{-2}$ in its vicinity.
@@
What are the consequences of the resonant non-equilibrium state? The two pumped quadratures correspond to $\sigma_{y}\tau_{x}$ – the easy direction for this mode (KIVC order), and $\sigma_{z}$ – the hard direction (sublattice polarization). The ratio of easy-to-hard quadratures is $\frac{1}{r}=\frac{{\cal E}_{\bf 0}^{\rm K}}{2\Delta_{\rm TK}}$. Remarkably, a phase which is otherwise inaccessible in an equilibrium setting, the KIVC analog in our model, plays a key role in the dynamics of the driven state. Moreover, this state has enhanced staggered sublattice correlations, which oscillate at half the driving frequency, $\omega_{d}/2$. The resultant phase of Floquet matter, enabled by coherent manipulation of this strongly interacting system, facilitates non-linear optical responses (e.g., second-harmonic generation [128, 129] due to the time-dependent inversion symmetry breaking) activated by pumping the system at a narrow frequency window.
IV.2.2 Sublattice Polarized Order (Fig. 4c)
@@
If the substrate-induced inversion symmetry breaking $\Delta_{\sigma}$ term is large enough, the preferred order parameter for the system is ${\cal O}_{\sigma}=\sigma_{z}$, i.e., sublattice polarization. The relevant ordered state hierarchy for this case is shown in Fig. 7a. The generators which keep the system in the low energy manifold of Eq. (55) are thus $\left\{{\cal M}_{i}\right\}=\sigma_{x}\tau_{x},\,\sigma_{y}\tau_{x},\,\sigma_{x}\tau_{y},\,\sigma_{y}\tau_{y}$.
@@
We may again organize then in canonically conjugate pairs, $\sigma_{x}\tau_{x}$ with $\sigma_{y}\tau_{x}$, and $\sigma_{x}\tau_{y}$ with $\sigma_{y}\tau_{y}$. The generator containing $\sigma_{y}$ in each pair generates a finite TIVC component in $|\Psi_{\phi}\rangle$, whereas the $\sigma_{x}$ generator promotes KIVC correlations (see Fig. 7b). It is thus natural to expect that the corresponding collective modes will be squeezed iff $\Delta_{\rm TK}\neq 0$. The lowest-lying modes corresponding to the two generator pairs are degenerate ($\tau_{x}$ and $\tau_{y}$ are on equal footing) and gapped (no continuous symmetry of the Hamiltonian is spontaneously broken by the ordered state).  [^8] We note that the $\Delta_{\sigma}$ scenario described here can be mapped onto the antiferromagnet discussed in Fig. 3c, with $\Delta_{\rm\sigma}$ playing the role of a staggered external magnetic field. To see that this is the case, simply perform a Pauli matrix replacement which conserves all commutation and anti commutation relations ($s^{i}$ and $p^{i}$ are Pauli matrices describing fictitious “spin” and “sublattice” degrees of freedom): The antiferromagnetic order is now $\sigma_{z}\to s^{z}p^{z}$, Néel vector rotations (or T-order generators) are $\sigma_{y}\tau_{x}\to s^{x}$, $\sigma_{y}\tau_{y}\to s^{y}$, and ferromagnetic ordering (or K-order generators) is given by generators $\sigma_{x}\tau_{x}\to s^{x}p^{z}$, $\sigma_{x}\tau_{y}\to s^{y}p^{z}$.
@@
As expected, the squeezing of lowest-lying collective modes is given by the breaking of T/K symmetry in the model. The modes have the expected gap at ${\bf Q}=0$,
${\cal E}^{\left(1,2\right)}_{\bf 0}=\sqrt{\left(\Delta_{\sigma}-\frac{u_{\zeta}}{2}\right)^{2}-\left(\frac{\Delta_{{\rm TK}}}{2}\right)^{2}}.$
The gap closes when the $\sigma_{z}$ order becomes exactly degenerate with the TIVC phase, i.e., when $\Delta_{\sigma}=\frac{u_{\zeta}+\Delta_{\rm TK}}{2}$.
The vicinity to a phase transition affects the parametric susceptibility dramatically. Consider small changes in the Hamiltonian parameters that affect the quantum geometry of electrons in the projected bands, modifying slightly the quantum metric length $\zeta$. In realistic devices, this can be the result of modulating gate voltages, electric displacement fields, or strain. More intriguing possibilities include modulation of external driving in Floquet systems, where the Floquet driving subtly controls single-particle properties [83]. The parametric efficiency to such a modification may be extracted by employing Eq. (40) (differentiating with respect to $\zeta$) and plugging it into Eq. (8),
@@
Notice that the left term in the denominator vanishes at the phase transition discussed above, resulting in the divergence of the efficiency, prominently shown in Fig. 7c–d. Once more, this can be understood as a consequence of the diverging fidelity susceptibility at the transition.
The dependence on $\zeta$ can be traced back to $u_{\zeta}$, and is maximized for $\zeta\approx 0.5$. This particular value is not universal – the maximizing value will typically depend on the range of interactions. Under our present assumptions, the $u$ interactions are maximally short-ranged, thus a change of length scale associated with the projected band electrons, i.e., $\sim\zeta$, beyond a certain value has limited effect. Generalizing this observation, small changes in quantum geometrical parameters are expected to have the most pronounce impact when the length-scale associated with the interaction is comparable to the one derived from the quantum metric. The interplay between these energy scales has been demonstrated to be crucial in the context of fractional Chern insulators [131, 132]. A parametric response probe may thus shed much needed light on the prospect of realizing these phases in possible candidate materials.
Figure 7: Pseudospin (sublattice) polarized order and its parametric driving (Sec. IV.2.2). (a) Energy hierarchy with a finite external sublattice potential $\Delta_{\sigma}$ in Eq. (53)]. The energy $u_{\zeta}$ brings the T/K orders down in energy, and $\Delta_{\rm TK}$ once again favors T-order over K. (b) The generalized Bloch sphere for the $\sigma$-order case, positioned the north pole. The two fluctuation quadratures correspond to the easy T-order, and the hard K-order, distinguished by the interaction energy $\Delta_{\rm TK}$. The anisotropy of the fluctuation cloud corresponds again to the vacuum. (c) Left: Parametric efficiency with respect to the quantum geometric parameter $\zeta$, Eq. (65). This parameter roughly corresponds to the Wannier spread in real-space of the wavefunctions corresponding to the active flat-bands (right schematic). The dashed line marks the phase boundary between the $\sigma$ phase below and the T-order stabilized above (hatched region), in accordance with the hierarchy in (a). Red line traces where $\eta_{{\rm par}.}^{\left(\zeta\right)}=1$. The parametric response is maximal when the interaction energy strongly depends on quantum geometry. The susceptibility diverges as one approaches the phase boundary, even at $\zeta$ much larger than its optimal value. Right: A vertical cut of the left colormap at $\zeta=0.65$, showing the divergence as the phase boundary is approached. Here, $\Delta_{\sigma}=u$.
@@
The response to driving in this scenario would be an oscillatory appearance of T-order and K-order components in the transient order parameter, with the amplitude of the former being stronger by the squeezing factor of $r^{-1}=\sqrt{1+\frac{\Delta_{{\rm TK}}}{\Delta_{\sigma}-\frac{u_{\zeta}+\Delta_{{\rm TK}}}{2}}}$ compared to the latter. We note that in the analogous twisted graphene system, the T-order, i.e., TIVC, couples to a real-space charge density with a Kekulé pattern [123, 134, 118, 82]. Thus, the parametrically driven state will give rise to a charge density wave oscillating at half the drive frequency, detectable using available time-resolved STM techniques [135, 136].